# When Parts Are Greater Than Sums: Individual LLM Components Can Outperform Full Models

### v2

@@ Abstract
This paper studies in-context learning (ICL) by decomposing the output of large language models into the individual contributions of attention heads and MLPs (components). We observe curious components: good-performing ones that individually do well on a classification task, even when the model performs poorly; bad-performing ones that do much worse than chance; and label-biased components that always predict the same label. We find that component accuracies are well-correlated across different demonstration sets and perturbations of prompt templates, even when the full-model accuracy varies greatly. Based on our findings, we propose component reweighting, which learns to linearly re-scale the component activations from a few labeled examples. Given $24$ labeled examples, our method improves by an average of $6.0\%$ accuracy points over $24$-shot ICL across 8 tasks on Llama-2-7B. Overall, this paper both enriches our understanding of ICL and provides a practical method for improvement by examining model internals.
@@ 1 Introduction
The rapid progress in large language models (LLMs) has popularized prompting, which guides LLMs to perform tasks with instructions or examples. Notably, in-context learning (ICL; Brown et al., 2020) adapts LLMs to a new task using only a few labeled examples without parameter updates. However, how LLMs react to the in-context examples is sometimes unintuitive Min et al. (2022b). Recently, Sclar et al. (2024) find that even for LLMs with instruction tuning Ouyang et al. (2022) or large size, adding a space or newline in prompts can greatly affect accuracy.
@@
We look into the LLM internals to understand what causes the surprising behavior across various ICL settings. Our work stands in contrast to prior studies on ICL, which often treat LLMs as black boxes and alter either the input Chen et al. (2023); Bertsch et al. (2024) or output Zhao et al. (2021); Holtzman et al. (2021). We introduce a new view of ICL by decomposing the output of an LLM into the sum of individual contributions of MLPs and attention heads, denoted “components.” Figure 1 reveals three types of curious components: good-performing (blue) that individually perform well or even outperform the full model, bad-performing (red) that perform below chance, and label-biased (green) that predict the same label on the entire test set. We observe these three types on Llama-2-7B, Llama-2-13B Touvron et al. (2023), and Mistral-Instruct-7B Jiang et al. (2023) across 8 tasks.
@@
We study the sensitivity of LLM components to multiple prompts, formed by different demonstrations and templates. We also construct contrast sets of templates—pairs of similar templates that yield large differences in ICL accuracy. Despite large variance in full-model accuracy, we find that component accuracies correlate well across different demonstrations ($r=0.80$ on average) and contrast set templates ($r=0.57$). The top-performing components in contrast set pairs overlap and achieve decent accuracy even when the full model performs near random (Figure 2). In contrast, the component accuracies of two sampled templates are less correlated ($r=0.34$). Further, good-performing components generalize well to out-of-distribution test sets. For instance, the top-1 component for MNLI outperforms the full Llama-2-13B model by $9.1\%$ on MedNLI; Figure 1 also shows that components are transferrable from SST2 to Yelp. We conclude that components are relatively consistent in their behavior across prompts and datasets.
@@
Inspired by our findings, we propose component reweighting. Compared to prior work that selects prompts to improve ICL accuracy Liu et al. (2022b), component reweighting softly selects components by learning weights from the same few-shot examples to scale component activations. Training these weights only involves learning a linear layer—less than a minute on one CPU. Overall, component reweighting better utilizes the same labeled examples, improving over $24$-shot ICL by $6.0\%,2.2\%,5.1\%$ on Llama-2-7B, Llama-2-13B, and Mistral-Instruct-7B, respectively, while enjoying similar inference speed as 4-shot ICL.
@@
Finally, we study the training dynamics of components using the Pythia pretraining checkpoints Biderman et al. (2023). During pretraining, good-performing components emerge well before the full model performs well. These findings suggest that LLMs acquire the internal ability to perform ICL early in training, but this ability only surfaces in the full model’s behavior later on. Our work conducts extensive analysis of LLM internals, which motivates a practical method to improve ICL; we hope to inspire future work that further sheds light on LLM internals in order to improve performance.
@@ 2.1 A New View of In-Context Learning
The residual stream of an LLM directly carries the information of the initial hidden state, every attention head, and every MLP, collectively named “components,” towards the output layer. We view this information as the direct contributions [^1] In comparison, a component has indirect contributions to the output by affecting other components in later layers. of components to the output logits, and derive a formula for logits, $\sum_{j}\mathbf{g}_{j}$, where $\mathbf{g}_{j}$ is the direct contribution of the component indexed by $j$. We can obtain the predictions of component $j$ with ${\arg\max}_{y\in\mathcal{Y}}\;\mathbf{g}_{j}$, and then calculate its individual ICL accuracy. Specifically, we derive $\mathbf{g}_{j}=U\cdot C_{j}$ in Eq. 8 below, where $U$ is the output embedding matrix and $C_{j}$ is the post-layernorm activations of component $j$. We name the operation $(C_{j}\mapsto U\cdot C_{j})$ as early decode, sharing the same spirit as nostalgebraist (2020); Geva et al. (2022), which interpret hidden representations by decoding through $U$. Compared to the standard view, we can directly study the behavior of individual components (Figure 2), characterizing them and scaling their contributions to the model output.
Figure 2: Left: Transformer decomposition. The components, MLPs and attention heads, are filled with blue and the blue lines show the flow of early decoding. Right: We can calculate the individual accuracy of every component after decomposition. Although a pair of templates that only differ slightly yield contrasting accuracies ($0.39$ vs. $0.89$ on AGNews with Llama-2-7B), the accuracies of their internal components are highly correlated. The top components for Template 1 overlap with the ones for Template 2 and achieve $>0.7$ accuracy despite the poor full-model accuracy.
@@ 2.2 A Walkthrough of the Decomposition
Mistral 7B Full $90.1_{\hskip 1.42271pt2.9}$ $\textbf{81.3}_{\hskip 1.42271pt2.1}$ $70.9_{\hskip 1.42271pt7.2}$ $58.5_{\hskip 1.42271pt4.2}$ $80.5_{\hskip 1.42271pt1.7}$ $56.1_{\hskip 1.42271pt5.0}$ $83.0_{\hskip 1.42271pt5.7}$ $\textbf{79.8}_{\hskip 1.42271pt1.4}$ $75.0$
@@ 3 Characterizing Components for ICL
We conduct in-context learning across 8 classification tasks on 3 LLMs: Llama-2-7B, Llama-2-13B, and Mistral-Instruct-7B. ICL is sensitive to prompts, so we randomly sample 5 disjoint sets of demonstrations formatted with 3 templates and report the standard deviation across the 15 runs. To avoid majority and recency biases Zhao et al. (2021), each prompt consists of the same number of demonstrations from every class in shuffled order. We use $K=3$ demonstrations for 3-way classification tasks and $K=4$ for the other tasks. Except for section 5.1, we refer to $K=\{3,4\}$ without further notice. We sample 2000 examples with balanced labels as the test set for every task. Please see A.1 for details about the tasks and templates.
@@ 3.1 Good and Bad-Performing Components
Across all the tasks and LLMs, we observe good-performing components that perform well or even outperform the full model and bad-performing components, which individually perform much worse than chance (blue and red dots in Figure 1, respectively). Table 1 compares the full model (Full) with the top-1 (Oracle-T1) and bottom-1 (Oracle-B1) components selected on the test set. On average, Oracle-T1 outperforms Full by $5.4\%,4.5\%,1.7\%$ on Llama-2-7B, Llama-2-13B, and Mistral-Instruct-7B, respectively; Oracle-B1 underperforms random guessing (Random) by $17.2\%,20.7\%,25.0\%$, respectively.
@@ 3.2 Label-Biased Components
Besides good and bad-performing components, we also observe label-biased components, which predict a certain label on the entire test set (the green dots in Figure 1). These components exist in all the tasks and LLMs we study, accounting for $29.1\%,26.4\%,22.8\%$ of components on average in Llama-2-7B, Llama-2-13B, and Mistral-Instruct-7B, respectively (see Table 7). When we prompt the model with all demonstrations of positive labels, a biased component still insists on predicting “negative” on the entire test set, and vice versa (A.2).
4 Transferability of Components
4.1 Transfer across Prompt Variants
Table 2 summarizes the results on Llama-2-7B. A.6 shows similar findings on other models. (1) The accuracies of the internal components are highly consistent across different choices of demonstrations, having strong correlations and an average of $0.54$ IoU. (2) The components have much weaker agreement across randomly sampled templates, having a near 0 IoU on BoolQ and QQP. (3) Nevertheless, there is agreement between minimally contrastive templates (Cst T), with an average correlation of $0.57$ across tasks, despite contrasting full-model accuracy. For example, Figure 2 demonstrates that full-model accuracy changes dramatically ($39\%$ vs $89\%$) in a minimal pair of templates, but internal components have a high correlation of $0.81$ and the pair shares top-performing components. Combining (2) and (3) suggests components behave similarly on similar templates, but this similarity decreases as the templates diverge.
@@ 4.2 Transfer to Out-of-Distribution Test Sets
Table 3 shows that Transfer-1 performs very close to Oracle-1 overall, suggesting that the top-performing components are consistent over data distribution. Moreover, Transfer-1 sometimes substantially outperforms Full, especially on MedNLI (random: $33.3\%$), suggesting that when LLMs perform poorly in ICL, it is likely that some internal components can do much better.
@@
Mistral 7B Full $\textbf{97.0}_{\hskip 1.42271pt0.5}$ $57.3_{\hskip 1.42271pt5.7}$ $\textbf{74.6}_{\hskip 1.42271pt3.5}$
@@
4.3 Transfer between Two Opposite Tasks
We conduct a case study of component transferability across instructions using Task069 and Task070 of Super-NaturalInstructions Wang et al. (2022b). The instruction for Task069 asks for correct answers, while Task070 asks for incorrect ones (“pick the one that makes less sense;” Figure 8). The examples for the two tasks are not parallel.
We find that Mistral-Instruct-7B achieves good accuracy across 15 runs on Task069 ($76.8\pm 2.4$), but below chance on Task070 ($40.6\pm 5.4$). We observe a strong negative correlation ($-0.60$ on average) between accuracies of every pair of runs from the two tasks. The worst-performing components in Task069 become the top-performing in Task070 and vice versa. The correlation suggests that the model has the ability to solve Task070, but misunderstands negation. Thus, we apply the Transfer-1 method (section 4.2) but select the worst-performing component from Task069 and then calculate its individual accuracy on Task070, achieving $58.7\pm 4.8$ accuracy across the 15 runs, an improvement of $\mathbf{18.1\%}$ over the full model. These results suggest that components behave consistently even across tasks with opposite instructions.

### v3

@@ Abstract
This paper studies in-context learning by decomposing the output of large language models into the individual contributions of attention heads and MLPs (components). We observe curious components: good-performing ones that individually do well on a classification task, even when the full model performs poorly; bad-performing ones that do much worse than chance; and label-biased components that always predict the same label. We find that component accuracies are well-correlated across different demonstration sets and perturbations of prompt templates. Based on our findings, we propose component reweighting, which learns to linearly re-scale the component activations from a few labeled examples. Given $24$ labeled examples, our method improves by an average of $6.0\%$ accuracy points over $24$-shot ICL across 8 tasks on Llama-2-7B. Overall, this paper both enriches our understanding of ICL and provides a practical method for improvement by examining model internals.
@@ 1 Introduction
The rapid progress in large language models (LLMs) has popularized prompting, which guides LLMs to perform tasks with instructions or examples. Notably, in-context learning (ICL; Brown et al., 2020) adapts LLMs to a new task using only a few labeled examples without parameter updates. However, how LLMs react to the in-context examples is sometimes unintuitive Min et al. (2022b). Recently, Sclar et al. (2024) and Voronov et al. (2024) find that even for instruction-tuned (Ouyang et al., 2022) or very large models, adding a space or newline in prompts can greatly affect accuracy.
@@
We look into the LLM internals to understand what causes the surprising behavior across various ICL settings. Our work stands in contrast to prior studies, which often treat LLMs as black boxes and alter either the input Chen et al. (2023); Bertsch et al. (2024) or output Zhao et al. (2021); Holtzman et al. (2021). We introduce a new view of ICL by decomposing the output of an LLM into the sum of individual contributions of MLPs and attention heads, denoted “components.” Figure 1 reveals three types of curious components: good-performing ones (blue) that individually perform well or even outperform the full model, bad-performing ones (red) that perform below chance, and label-biased ones (green) that predict the same label on the entire test set. We observe these three classes of components on Llama-2-7B, Llama-2-13B Touvron et al. (2023), Llama-3-8B Dubey et al. (2024), and Mistral-Instruct-7B Jiang et al. (2023) across 8 classification tasks.
@@
We study the sensitivity of LLM components to multiple prompts formed by different demonstrations and templates. We also construct contrast sets of templates—pairs of similar templates that yield large differences in ICL accuracy. Despite large variance in full-model accuracy, we find that component accuracies correlate well across different demonstrations ($r=0.80$ on average) and contrast set templates ($r=0.57$). The top-performing components in contrast set pairs overlap and achieve decent accuracy even when the full model performs near random (Figure 2). Nonetheless, the component accuracies of two sampled templates are less correlated ($r=0.34$). Further, good-performing components generalize well to out-of-distribution test sets. For instance, the top-1 component for MNLI outperforms the full Llama-2-13B model by $9.1\%$ on MedNLI; Figure 1 also shows that components are transferrable from SST2 to Yelp. We conclude that components are relatively consistent in their behavior across prompts and datasets.
@@
Inspired by our findings, we propose component reweighting. Compared to prior work that selects prompts from a large pool of labeled data to improve ICL accuracy Liu et al. (2022b), component reweighting softly selects components by learning weights from few-shot examples to scale component activations. Training these weights only involves learning a linear layer, which takes less than a minute on one CPU. Overall, component reweighting better utilizes the same labeled examples, improving over $24$-shot ICL by $6.0\%,2.2\%,5.1\%,1.6\%$ on Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively. At the same time, it enjoys similar inference speed as 4-shot ICL.
@@
Finally, we study the training dynamics of components using the Pythia pretraining checkpoints Biderman et al. (2023). During pretraining, good-performing components emerge well before the full model performs well. These findings suggest that LLMs acquire the internal ability to perform ICL early in training, but this ability only surfaces in the full model’s behavior later on.
Overall, our work conducts extensive analysis of LLM internals, which motivates a practical method to improve ICL. We hope to inspire future work that further sheds light on LLM internals in order to improve performance. Our implementation is available at https://github.com/terarachang/LLMDecomp.
@@ 2.1 A New View of In-Context Learning
The residual stream of an LLM directly carries the information of the initial hidden state, every attention head, and every MLP, collectively named “components,” towards the output layer. We view this information as the direct contributions [^1] In comparison, a component has indirect contributions to the output by affecting other components in later layers Wang et al. (2023a). This paper focuses on direction contributions. of components to the output logits, and derive a formula for logits, $\sum_{j}\mathbf{g}_{j}$, where $\mathbf{g}_{j}$ is the direct contribution of the component indexed by $j$. We can obtain the predictions of component $j$ with ${\arg\max}_{y\in\mathcal{Y}}\;\mathbf{g}_{j}$, and then calculate its individual ICL accuracy. Specifically, we derive $\mathbf{g}_{j}=U\cdot C_{j}$ in Eq. 8 below, where $U$ is the output embedding matrix and $C_{j}$ is the post-layernorm activations of component $j$. We name the operation $(C_{j}\mapsto U\cdot C_{j})$ as early decode, sharing the same spirit as nostalgebraist (2020) and Geva et al. (2022), which interpret hidden representations by decoding through $U$. Compared to the standard view, we can directly study the behavior of individual components (Figure 2), characterizing them and scaling their contributions to the model output.
Figure 2: Left: Transformer decomposition. The components—MLPs and attention heads—are filled with blue, and the blue lines show the flow of early decoding. Right: We can calculate the individual accuracy of every component after decomposition. Although a pair of templates that only differ slightly yield very different accuracies ($0.39$ vs. $0.89$ on AGNews with Llama-2-7B), the accuracies of their internal components are highly correlated. The top components for Template 1 overlap with the ones for Template 2 and achieve $>0.7$ accuracy despite the poor full-model accuracy.
@@ 2.2 A Walkthrough of the Decomposition
Mistral Ins 7B Full $90.1_{\hskip 1.42271pt2.9}$ $\textbf{81.3}_{\hskip 1.42271pt2.1}$ $70.9_{\hskip 1.42271pt7.2}$ $58.5_{\hskip 1.42271pt4.2}$ $80.5_{\hskip 1.42271pt1.7}$ $56.1_{\hskip 1.42271pt5.0}$ $83.0_{\hskip 1.42271pt5.7}$ $\textbf{79.8}_{\hskip 1.42271pt1.4}$ $75.0$
@@
Llama3 8B Full $91.4_{\hskip 1.42271pt1.7}$ $\textbf{79.2}_{\hskip 1.42271pt7.2}$ $74.0_{\hskip 1.42271pt8.0}$ $58.7_{\hskip 1.42271pt4.7}$ $\textbf{76.5}_{\hskip 1.42271pt2.2}$ $59.4_{\hskip 1.42271pt3.7}$ $\textbf{84.0}_{\hskip 1.42271pt6.6}$ $\textbf{87.4}_{\hskip 1.42271pt5.5}$ $76.3$
Oracle-T1 $\textbf{92.3}_{\hskip 1.42271pt1.0}$ $77.4_{\hskip 1.42271pt7.3}$ $\textbf{77.4}_{\hskip 1.42271pt3.7}$ $\textbf{64.5}_{\hskip 1.42271pt2.7}$ $76.3_{\hskip 1.42271pt2.9}$ $\textbf{60.7}_{\hskip 1.42271pt1.5}$ $81.4_{\hskip 1.42271pt5.7}$ $86.0_{\hskip 1.42271pt5.9}$ 77.0
Oracle-B1 $\phantom{1}9.0_{\hskip 1.42271pt0.9}$ $22.5_{\hskip 1.42271pt7.4}$ $23.5_{\hskip 1.42271pt3.9}$ $36.7_{\hskip 1.42271pt3.3}$ $23.4_{\hskip 1.42271pt2.1}$ $10.0_{\hskip 1.42271pt4.2}$ $\phantom{1}1.4_{\hskip 1.42271pt0.6}$ $\phantom{1}1.9_{\hskip 1.42271pt0.9}$ $16.0$
@@ 3 Characterizing Components for ICL
We conduct in-context learning across 8 classification tasks on 4 LLMs: Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B. ICL is sensitive to prompts, so we randomly sample 5 disjoint sets of demonstrations formatted with 3 templates and report the standard deviation across the 15 runs. To avoid majority and recency biases Zhao et al. (2021), each prompt consists of the same number of demonstrations from every class in shuffled order. We use $K=3$ demonstrations for 3-way classification tasks and $K=4$ for the other tasks. Except for section 5.1, we refer to $K=\{3,4\}$ without further notice. We sample 2000 examples with balanced labels as the test set for every task. Please see A.1 for details about the tasks and templates.
@@ 3.1 Good and Bad-Performing Components
Across all the tasks and LLMs, we observe good-performing components that perform well or even outperform the full model, and bad-performing components that individually perform much worse than chance (blue and red dots in Figure 1, respectively). Table 1 compares the full model (Full) with the top-1 (Oracle-T1) and bottom-1 (Oracle-B1) components selected on the test set. On average, Oracle-T1 outperforms Full by $5.4\%,4.5\%,1.7\%,0.7\%$ on Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively; Oracle-B1 underperforms random guessing (Random) by $17.2\%,20.7\%,25.0\%,25.7\%$.
@@ 3.2 Label-Biased Components
Besides good and bad-performing components, we also observe label-biased components, which predict a certain label on the entire test set (the green dots in Figure 1). These components exist in all the tasks and LLMs we study, accounting for $29.1\%,26.4\%,22.8\%,29.7\%$ of components on average in Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively (Table 5). In A.2, we show that even when we prompt the model with all demonstrations of positive labels, the most biased component still insists on predicting “negative” on the entire test set, and vice versa.
3.3 Mechanistic Understanding of Bad-Performing Heads
Prior work studies the mechanism of certain components in LLMs, showing that there are negative mover attention heads that write in the opposite direction of the expected answer Wang et al. (2023a) and copy suppression heads that suppress the prediction of a prior token in the context McDougall et al. (2023). Inspired by them, we investigate the mechanism behind bad-performing heads identified by our decomposition. We focus on label tokens in the context, as Wang et al. (2023b) show that label words serve as anchors in ICL.
We conduct a case study on Llama-2-7B with 4-shot balanced in-context examples from SST2. We examine the bottom-5 attention heads that have the worst ICL accuracy on SST2. We find that three of these heads, L19H15, L15H14, and L18H9, assign top attention probabilities to all 4 label tokens of the 4-shot in-context examples when predicting test examples. Furthermore, despite their poor ICL accuracy, these heads actually assign higher attention to the correct in-context label tokens than the incorrect ones most of the time ($>70\%$ of the test examples). In other words, when a test example has a positive label, these heads assign higher attention [^3] We average the attention probabilities of the same label tokens and then compare the average ones of the two labels. to the tokens “positive” in the context than the tokens “negative”. We also observe that the more the heads attend to “positive” in the context, the lower the inner product between the head and the output embedding of the token “positive”, with the correlation $r=-0.97,-0.96,-0.89$ for L19H15, L15H14, and L18H9, respectively.
[…]
@@ 3.3 Mechanistic Understanding of Bad-Performing Heads
4 Transferability of Components
4.1 Transfer across Prompt Variants
Table 2 summarizes the results on Llama-2-7B; A.6 shows similar findings on other models. (1) The accuracies of the internal components are highly consistent across different choices of demonstrations, having strong correlations and an average of $0.54$ IoU. (2) The components have much weaker agreement across randomly sampled templates, having a near 0 IoU on BoolQ and QQP. (3) Nevertheless, there is agreement between minimally contrastive templates (Cst T), with an average correlation of $0.57$ across tasks, despite contrasting full-model accuracy. For example, Figure 2 demonstrates that full-model accuracy changes dramatically ($39\%$ vs $89\%$) in a minimal pair of templates, but internal components have a high correlation of $0.81$ and the pair shares top-performing components. Combining (2) and (3) suggests components behave similarly on similar templates, but this similarity decreases as the templates diverge.
@@ 4.2 Transfer to Out-of-Distribution Test Sets
Table 3 shows that Transfer-1 closely matches Oracle-1 overall, suggesting that the top-performing components are transferable across data distribution. Moreover, Transfer-1 sometimes outperforms Full, especially on Llama2 models, showing the hidden abilities of the internal components.
4.3 Transfer between Two Opposite Tasks
We conduct a case study of component transferability across instructions using Task069 and Task070 of Super-NaturalInstructions Wang et al. (2022b), both of which are binary abductive NLI tasks Bhagavatula et al. (2020). The instruction for Task069 asks for correct answers, while Task070 asks for incorrect ones (“pick the one that makes less sense;” see Figure 7 for the full instructions). Examples in the two tasks are not parallel.
We find that Mistral-Instruct-7B achieves good accuracy across 15 runs on Task069 ($76.8\pm 2.4$), but below chance on Task070 ($40.6\pm 5.4$). We observe a strong negative correlation, $r=-0.60$ on average, between the component accuracies of the two tasks. The worst-performing components in Task069 become the top-performing in Task070 and vice versa. The correlation suggests that the model has the ability to solve Task070, but misunderstands negation. Thus, we apply the Transfer-1 method (section 4.2) but select the worst-performing component from Task069 and then calculate its individual accuracy on Task070. Transfer-1 achieves $58.7\pm 4.8$ accuracy across the 15 runs, an improvement of $\mathbf{18.1\%}$ over the full model. These results suggest that components behave consistently even across tasks with opposite instructions, as the active components in Task069 are also active in Task070.
@@ 4.3 Transfer between Two Opposite Tasks
Mistral Ins 7B Full $\textbf{97.0}_{\hskip 1.42271pt0.5}$ $57.3_{\hskip 1.42271pt5.7}$ $\textbf{74.6}_{\hskip 1.42271pt3.5}$
@@
Llama3 8B Full $\textbf{97.8}_{\hskip 1.42271pt0.4}$ $61.0_{\hskip 1.42271pt2.2}$ $\textbf{77.3}_{\hskip 1.42271pt7.5}$
Transfer-1 $95.9_{\hskip 1.42271pt4.4}$ $\textbf{61.3}_{\hskip 1.42271pt0.8}$ $73.9_{\hskip 1.42271pt8.4}$
Oracle-1 $97.9_{\hskip 1.42271pt0.5}$ $61.6_{\hskip 1.42271pt0.6}$ $74.8_{\hskip 1.42271pt8.9}$